# Magnetized neutral 2SC color superconductivity and possible origin of the inner magnetic field of magnetars

### v1

@@ Abstract
In this paper the neutral 2SC phase of color superconductivity is investigated in the presence of a magnetic field and for diquark coupling constants and baryonic densities that are expected to characterize neutron stars. Specifically, the behavior of the charged gluons Meissner masses is investigated in the parameter region of interest taking into account in addition the contribution of a rotated magnetic field. It is found that up to moderately-high diquark coupling constants the mentioned Meissner masses become tachyonic independently of the applied magnetic field amplitude, hence signalizing the chromomagnetic instability of this phase. To remove the instability, it is proposed the restructuring of the system ground state, which now will be formed by vortices of the rotated charged gluons. These vortices boost the applied magnetic field having the most significant increase for relatively low applied magnetic fields. Finally, considering that with the stellar rotational frequency observed for magnetars a field of order $10^{8}$ G can be generated by dynamo effect. Then, we show that by the boosting effect just described the field can be amplified to $10^{17}$ G that is in the range of inner core fields expected for magnetars. Thus, we conclude that the described mechanism could be the one responsible for the large fields characterizing magnetars if the core of this compact objects are formed by neutral 2SC matter.
@@ I Introduction
It is of great interest to the nuclear physics and astrophysics communities to elucidate what are the physical characteristics of the super-dense matter that forms the interior of neutron stars (NS). In this context, to understand the effect of a strong magnetic field on that medium is essential due to the fact that magnetic fields of different intensities populate the entire landscape of these compact objects. The surface magnetic fields of some radio pulsars are of order $10^{8}-10^{12}$ G NS-B-1 . There are even some special compact objects called magnetars NS-B-2 whose surface magnetic fields are of order $10^{14}-10^{15}$ G. Those values have been inferred from spectroscopic and spin-down studies of soft-gamma ray repeaters (SGRs) and anomalous x-ray pulsars (AXPs). In addition, we should take into account that the inner core magnetic fields of magnetars can be even larger, which follows from the magnetic field flux conservation in stellar media possessing very large electric conductivities. By using different arguments the inner fields have been estimated to range from $10^{17}$ G to $10^{20}$ G magnetizedfermions ; Theor-estimations . The fact that the stellar inner medium will be under the effect of such strong magnetic fields, has motivated many works focused on the study of the equation of state (EOS) of magnetized NS (see insignificance and references there). On the other hand, to understand what originates the tremendous magnetic fields that characterizes magnetars remains as a question under investigation.
@@
Quark matter under sufficiently cold and dense conditions is expected to be realized in the interior of massive NS Weber ; Quark-stars . Under such conditions, the existence of a color superconducting (CS) phase is unavoidable. This phase is characterized by the formation of Cooper pairs of quarks due to the attractive interaction originating either from the perturbative one-gluon exchange in color antisymmetric channel at ultra high baryon density or from the instanton effects at moderate density. When the masses of the $u$, $d$ and $s$ quarks can be ignored compared to the baryon chemical potential and all three flavors take part in the pairing, the ground state is characterized by a spin-zero CS condensate giving rise to what is called as the Color-Flavor-Locked (CFL) phase CFL . This phase is the well-established ground state at asymptotic high density and low temperature. On the other hand, at moderate densities, where it is expected that the quark interaction is stronger, if the density is high enough to release quarks from inside the hadrons and sufficiently low to decouple the $s$ quark dynamics, the pairing involves only two light flavors and the ground state is a two-flavor supreconducting (2SC) phase Igor-Rischke .
The strength of the existing strong magnetic fields in the interior of NS can create the conditions for observable QCD effects B_HIC-2 . These effects can be prominent because the generated magnetic fields are of the order of or higher than the QCD scale, $eB>\Lambda^{2}_{QCD}$. Specifically, in CS at asymptotically high densities the CFL phase is split in three new phases depending on the magnetic field strength Phases ; E-V-Review . The physical scales with which the magnetic field strength has to compete to produce the new phases are the CFL gap, the rotationally charged gluon Meissner mass and the baryonic chemical potential. A sufficiently strong magnetic field, for example, changes the symmetry of the CFL phase modifying the pattern of the diquark condensates MCFL ; Bo , hence defining a new phase, which is called the magnetic CFL phase (MCFL).
@@
There is another physical factor affecting the stellar medium: the electric neutrality condition of the medium (given by $\partial\Omega/\partial\mu_{e}=0$, with $\Omega$ the system thermodynamic potential and $\mu_{e}$ the electric chemical potential), which can also affect the CS phase specially at moderate densities where it is expected that the so called 2CS phase can be realized. Some authors have already investigated the neutral 2SC phase in the presence of a magnetic field Jaikumar-1 -Mishra . Those works had been constrained to study the evolution of the quark gap $\Delta$ and constituent quark mass $m$ with increasing densities and magnetic fields. Nevertheless, the effect of an applied magnetic field on the gluon sector of such an electrically neutral system has been ignored up to now.
@@
Considering that quark matter in the interior of NS must be neutral with respect to both electric and color charges and in $\beta$-equilibrium, it was found Igor that the pairing dynamics within the 2SC phase results affected due to the mismatch between the Fermi spheres of quarks with different flavors. As a consequence, gapless modes in the spectrum of quasiparticles appear when the mismatch $\delta\mu=\mu_{e}/2$ satisfies the condition $10<\Delta/\delta\mu<$ 1. This unusual phase is called gapless 2SC phase (g2SC). Further calculations Huang revealed the other important consequence that the square of the Meissner mass of some gluons are negative signaling an instability, which was named the chromomagnetic instability. It is worth to emphasize that the gapless superconductivity itself is not the reason for such an instability since, for example, in the gapped 2SC phase the squared Meissner mass of some gluons are also negative when $0<\Delta/\delta\mu<\sqrt{2}$. The chromomagnetic instability poses a major challenge in the studies of CS phases with quark pairing mismatch. Although a consensual conclusion to remove such an instability hasn’t been reached, possible alternative solutions already exist in the literature. In this context, inhomogeneous solutions seem to play an important role. Some of them spontaneously break translational invariance Reddy -Hashimoto through inhomogeneous quark-quark condensates which involves a momentum-dependent gap. These kind of inhomogeneous CS phases are based on the idea of Larkin and Ovchinnikov Larkin and Fulde and Ferrell Ferrell originally applied to condensed matter. Another inhomogeneous phase that breaks rotational invariance has been also considered where gluon vortices are induced together with the spontaneous generation of a magnetic field Ferrer-Vortices . This, second scenario is motivated by the solution to the so-called ”zero-mode problem” that takes place for charged spin-one field in the presence of a magnetic field larger than the field square Meissner mass Yang-Mills -String Theory .
@@
In the present work, we will show that in the presence of a magnetic field the unstable neutral 2SC phase will be stabilized by restructuring its ground state by the formation of a gluon vortex condensate as it follows from the well known solution of the ”zero-mode problem,” which has been already found in different contexts as for Yang-Mills fields in Yang-Mills , for the $W^{\pm}_{\mu}$ bosons of the electroweak theory in W-EW , for the charged gluons of the MCFL phase of CS in MCFL-Vortices and even for higher-spin fields in the context of string theory in String Theory . In the magnetized neutral 2SC system the gluon vortex solution will be even reinforced for certain density regions in the presence of a magnetic field as we will show in this work.
@@
Something else to notice is that an important characteristic of spin-zero CS phases is the lack of Meissiner effect in contrast to the conventional superconductivity. Although the original electromagnetic $U(1)_{em}$ symmetry is broken by the formation of quark Cooper pairs, a residual $\widetilde{U}(1)$ symmetry still remains. The massless gauge field associated with this symmetry in the 2SC phase is given by the linear combination of the conventional photon field $A_{\mu}$ and the 8th-gluon field $G^{8}_{\mu}$ Rotated Fields as $\widetilde{A}_{\mu}=\cos\theta A_{\mu}-\sin\theta G^{8}_{\mu}$. The field $\widetilde{A}_{\mu}$ plays the role of a long-range in-medium electromagnetic field, which is also called the rotated electromagnetic field, while the orthogonal combination $\widetilde{G}_{\mu}^{8}=\sin\theta A_{\mu}+\cos\theta G^{8}_{\mu}$ is massive. The mixing angle depends on the coupling constants through the relation $\tan\theta=e/\sqrt{3}g$, with $g$ and $e$ the strong and electromagnetic coupling constants, respectively. In this paper we will investigate the effect of the rotated magnetic field associated with $\widetilde{A}_{\mu}$ on neutral 2SC matter and specially on the Meissner masses of gluons that acquire rotated electric charges, $G^{(4,5,6,7)}_{\mu}$, and which are the ones associated with the tachyonic modes in this medium. The mechanism that serves to remove in this context the chromomagnetic instability is due to the formation of gluonic vortices together with the generation of a magnetic field. Thus, we will also show how much the applied magnetic field can be boosted by the generation of the gluon vortices that can be induced at the moderately high baryon density expected to be reached in NS. Finally, based on those results, one of the main outcomes of this paper will be to present a natural mechanism to explain the possible origin of the strong magnetic fields exhibited by magnetars.
@@
The rest of the paper is organized as follows: In Section 2, we review the NJL formulation of the 2SC neutral phase of CS in the presence of a uniform magnetic field. The results for the gap, $\Delta$, and electric chemical potential, $\mu_{e}$, found in this section will be used then in the subsequent sections. In Section 3, the Meissner mass of the rotationally charged gluons is calculated for different values of the diquark coupling constant in the range of baryonic densities that are considered to be of interest for NS. It will be shown that for coupling constants not extremely high, the Meissner mass exhibits a tachyonic behavior at moderately high magnetic fields in the whole rage of densities under consideration. This is what has been called the chromamegnetic instability phenomenon. It will be discussed how this instability is removed by redefining the system ground state through the formation of gluon vortices. A peculiar effect generated by these vortices is the amplification of the magnitude of the applied magnetic field. By numerical calculations we will see that small fields up to $10^{15}$ G will be boosted in two orders of magnitude. This effect will be used then in Section 4 as a mechanism to produce the high strength that the magnetar magnetic fields can reach in its interior. In Section 5 we present the concluding remarks.
@@ II Cooper pair condensate at ${\bm{B}}\neq 0$ in the neutral medium
We consider in this paper a two flavor NJL model of massless quarks in the presence of an external constant and uniform rotated magnetic field (From now on we are going to ignore the term ”rotated” but it will be indicated by a wavy-hat notation on top of the magnitudes). The Lagrangian reads
${\cal L}=\bar{\psi}\left[i\gamma^{\mu}(\partial_{\mu}+i\tilde{e}\tilde{Q}\tilde{A}_{\mu})+\mu\gamma^{0}\right]\psi+G_{S}\left[(\bar{\psi}\psi)^{2}+(\bar{\psi}i\gamma^{5}{\bf\tau}\psi)^{2}\right]+G_{D}\left(\bar{\psi}i\gamma^{5}T^{2}\tau^{2}\psi^{C}\right)\left(\bar{\psi}^{C}i\gamma^{5}T^{2}\tau^{2}\psi\right),$ (1)
where $\psi^{T}=(u,d)$ is the the quark doublet and $\psi^{C}=C\bar{\psi}^{T}$ is its charge conjugate with $C=i\gamma^{2}\gamma^{0}$ the charge conjugation operator, $G_{S},G_{D}$ are the coupling constants of the scalar and diquark channels of the four-fermion interaction theory respectively. Here, $T^{2}$ is the 2nd Gell-Mann (antisymmetric) matrix in color space and $\tau^{2}$ is the Pauli (antisymmetric) matrix in flavor space. $\tilde{Q}$ is the rotated charge matrix defined as
@@
in unite of $\tilde{e}=\frac{\sqrt{3}ge}{\sqrt{3g^{2}+e^{2}}}>0$. Here $Q_{f}={\rm diag}(2/3,-1/3)$ is the original electromagnetic charge matrix for quarks and $T_{c}^{8}={\rm diag}(1/\sqrt{3},1/\sqrt{3},-2/\sqrt{3})$ is the 8th Gell-Mann matrix. To be explicit, the $\tilde{Q}$ charge for different quarks are given in Table I.
@@
In this convention, the only diquark pairs are $u_{r}d_{g}$, $u_{g}d_{r}$. These pairs are neutral with respect to $\tilde{Q}$. This is the reason why there is no Meissner effect for $\tilde{B}$ due to the absence of a charged ground state.
@@
The quark chemical potentials related with electrical and color neutralities, as well as $\beta$-equilibrium can be expressed by the following matrix in color-flavor space
@@
The notations $\bar{\mu}=\mu-\frac{1}{6}\mu_{e}+\frac{1}{3}\mu_{8}$, $\delta\mu_{e}=\frac{1}{2}\mu_{e}$ and $\delta\mu_{8}=\mu_{8}$ have been used. The chemical potentials $\mu_{e}$ and $\mu_{8}$ are dynamical parameters that have to be determined from the electrical and color neutrality equations. The color chemical potential $\mu_{8}$ is usually very small compared with the other chemical potentials $\mu_{e}$ and $\mu$. We therefore will assume $\mu_{8}=0$ in the following calculations since as has been found by other authors it will not significantly change the results.
@@
which is obtained from Eq. (1) by assuming $\Delta=2G_{D}\langle{\bar{\psi}}^{C}i\gamma^{5}T^{2}\tau^{2}\psi\rangle$ and making the mean field approximation. Note that we are not considering the chiral condensation. From previous results Jaikumar-1 -Mishra , we see that for the density region of greatest interest ($\mu>300$ MeV) the chiral condensate already vanished.
@@
Figure 1: The solutions of the gap $\Delta$ and mismatch $\delta\mu_{e}$ versus baryonic chemical potential for different diquark couplings at a magnetic field value of $10^{16}$.
Figure 2: The solutions of the gap $\Delta$ and mismatch $\delta\mu_{e}$ versus baryonic chemical potential for different diquark couplings at a magnetic field value of $10^{18}$ G.
Thus we obtain the thermodynamical potential,
@@
with the contribution in the parenthesis coming from the electrons, which must be introduced to maintain the system neutrality and $\beta$ equilibrium.
Carrying out the sum in Mastubara frequencies, we finally have
@@
In Figs. 1 and 2, we present for two different values of magnetic fields the solutions of the minimum equations (36) in the range of baryonic chemical potentials from $250\ {\rm MeV}$ to $500\ {\rm MeV}$, which are expected to correspond to the typical density region of compact stars. There, we can observe that there is a jump in $\Delta$ from zero to a higher value at relatively low baryonic chemical potentials for lower values of the diquark coupling. This jump is signalizing the formation of the Fermi sphere for those parameter values. The jump in $\Delta$ is accompanied by a jump in the mismatch $\delta\mu_{e}$ showing how the formation of the Fermi sphere is affecting the value of the electric chemical potential at different diquark couplings. At very large coupling the formation of the Cooper pairs is present in the whole range of baryonic chemical potentials under consideration. Comparing Fig. 1 and 2 we notice that the magnetic field does not significantly affect the behavior of $\Delta$ and $\delta\mu_{e}$ in the parameter range under consideration. As expected, the coupling strength $\eta$ affects significantly the value of the gap $\Delta$ producing a significant increase as it is increased, the electric chemical potential, on the other hand, is also increased indicating that as more quarks condense in Cooper pairs having a positive conventional electric charge, the system needs a larger number of electrons to keep its electric neutrality.
Figure 3: The ratio of the squared Meissner mass of charged gluons to the magnetic mass for different diquark couplings at a magnetic field with value of $10^{16}G$.
Figure 4: The ratio of the squared Meissner mass of charged gluons to the magnetic mass for different diquark couplings at a magnetic field with value of $10^{18}G$.
@@
The gluon screening properties in the neutral 2SC phase can be studied by calculating the gluon polarization tensors considered in detail in Ref. Huang . In principle, for sufficiently high magnetic fields the polarization tensor of charged particles can change their Lorentz tensorial structure showing a space anisotropy determined by the field direction Pi_Anisotropy . Nevertheless, for astrophysical applications to compact objects since $eB\ll\mu^{2}$, the anisotropy in the Lorentz structure of the polarization operator becomes a second-order effect that can be neglected, and hence we can keep the polarization operator structure characteristic of a dense medium at zero field reducing the effect of the magnetic field to its dependence on $\Delta$ and $\mu_{e}$.
In Huang it was found that while the Meissner masses of the three gluons that correspond to the unbroken $SU(2)_{c}$ subgroup are vanishing as expected, the charged gluons that correspond to the broken generators of the $SU(3)_{c}$ group could be imaginary in some parameter regions. The explicit Meissner masses for the charged and the rotated $8$th gluon fields are given respectively by

### v2

@@ Abstract
In this paper the neutral 2SC phase of color superconductivity is investigated in the presence of a magnetic field and for diquark coupling constants and baryonic densities that are expected to characterize neutron stars. Specifically, the behavior of the charged gluons Meissner masses is investigated in the parameter region of interest taking into account, in addition, the contribution of a rotated magnetic field. It is found that up to moderately-high diquark coupling constants the mentioned Meissner masses become tachyonic independently of the applied magnetic-field amplitude, hence signalizing the chromomagnetic instability of this phase. To remove the instability, the restructuring of the system ground state is proposed, which now will be formed by vortices of the rotated charged gluons. These vortices boost the applied magnetic field having the most significant increase for relatively low applied magnetic fields. Finally, considering that with the stellar rotational frequency observed for magnetars a field of the order of $10^{8}$ G can be generated by dynamo effect, we show that by the boosting effect just described the field can be amplified to $10^{17}$ G that is in the range of inner core fields expected for magnetars. Thus, we conclude that the described mechanism could be the one responsible for the large fields characterizing magnetars if the cores of these compact objects are formed by neutral 2SC matter.
@@ I Introduction
It is of great interest to the nuclear physics and astrophysics communities to elucidate what are the physical characteristics of the superdense matter that forms the interior of neutron stars (NS). In this context, to understand the effect of a strong magnetic field on that medium is essential due to the fact that magnetic fields of different intensities populate the entire landscape of these compact objects. The surface magnetic fields of some radio pulsars are of the order of $10^{8}-10^{12}$ G NS-B-1 . There are even some special compact objects called magnetars NS-B-2 whose surface magnetic fields are of the order of $10^{14}-10^{15}$ G. Those values have been inferred from spectroscopic and spin-down studies of soft-gamma ray repeaters (SGRs) and anomalous x-ray pulsars (AXPs). In addition, we should take into account that the inner core magnetic fields of magnetars can be even larger, which follows from the magnetic-field flux conservation in stellar media possessing very large electric conductivities. By using different arguments the inner fields have been estimated to range from $10^{17}$ to $10^{20}$ G magnetizedfermions ; Theor-estimations . The fact that the stellar inner medium will be under the effect of such strong magnetic fields, has motivated many works focused on the study of the equation of state of magnetized NS (see insignificance and references there). On the other hand, to understand what originates the tremendous magnetic fields that characterizes magnetars remains as a question under investigation.
@@
Quark matter under sufficiently cold and dense conditions is expected to be realized in the interior of massive NS Weber ; Quark-stars . Under such conditions, the existence of a color superconducting (CS) phase is unavoidable. This phase is characterized by the formation of Cooper pairs of quarks due to the attractive interaction originating either from the perturbative one-gluon exchange in the color antisymmetric channel at ultrahigh baryon density or from the instanton effects at moderate density. When the masses of the $u$, $d$ and $s$ quarks can be ignored compared to the baryon chemical potential and all three flavors take part in the pairing, the ground state is characterized by a spin-zero CS condensate giving rise to what is called as the Color-Flavor-Locked (CFL) phase CFL . This phase is the well-established ground state at asymptotic high density and low temperature. On the other hand, at moderate densities, where it is expected that the quark interaction is stronger, if the density is high enough to release quarks from inside the hadrons and sufficiently low to decouple the $s$ quark dynamics, the pairing involves only two light flavors and the ground state is a two-flavor superconducting (2SC) phase Igor-Rischke .
@@
In investigating compact stars, there is another physical factor to be considered: the electric neutrality of the medium (given by $\partial\Omega/\partial\mu_{e}=0$, with $\Omega$ the system thermodynamic potential and $\mu_{e}$ the electric chemical potential), which can also affect the CS phase specially at moderate densities where it is expected that the so-called 2SC phase can be realized. Some authors have already investigated the neutral 2SC phase in the presence of a magnetic field Jaikumar-1 -Mishra . Those works had been constrained to study the evolution of the quark gap $\Delta$ and constituent quark mass $m$ with increasing densities and magnetic fields. Nevertheless, the effect of an applied magnetic field on the gluon sector of such an electrically neutral system has been ignored up to now.
@@
Considering that quark matter in the interior of NS must be neutral with respect to both electric and color charges and in $\beta$ equilibrium, (i.e., the weak beta-equilibrium condition $\mu_{u}+\mu_{e}-\mu_{\nu_{e}}=\mu_{d}$ should be satisfied, with $\mu_{u}$ the chemical potential of the $u$ quark, $\mu_{d}$ the one for the $d$ quark, $\mu_{e}$ the electron chemical potential and $\mu_{\nu_{e}}$ that of the electron neutrinos, which will be here ignored considering stars in the post neutrino-emission epoch.) it was found Igor that the pairing dynamics within the 2SC phase results affected due to the mismatch between the Fermi spheres of quarks with different flavors. As a consequence, gapless modes in the spectrum of quasiparticles appear when the mismatch $\delta\mu_{e}=\mu_{e}/2$ satisfies the condition $10<\Delta/\delta\mu_{e}<$ 1. This unusual phase is called the gapless 2SC phase (g2SC). Further calculations Huang revealed the other important consequence that the square of the Meissner mass of some gluons are negative signaling an instability, which was named the chromomagnetic instability. It is worth to emphasize that the gapless superconductivity itself is not the reason for such an instability since, for example, in the gapped 2SC phase the squared Meissner mass of some gluons are also negative when $1<\Delta/\delta\mu_{e}<\sqrt{2}$. The chromomagnetic instability poses a major challenge in the studies of CS phases with quark pairing mismatch. Although a consensual conclusion to remove such an instability hasn’t been reached, possible alternative solutions already exist in the literature. In this context, inhomogeneous solutions seem to play an important role. Some of them spontaneously break translational invariance Reddy -Hashimoto through inhomogeneous quark-quark condensates which involves a momentum-dependent gap. These kind of inhomogeneous CS phases are based on the idea of Larkin and Ovchinnikov Larkin and Fulde and Ferrell Ferrell originally applied to condensed matter. Another inhomogeneous phase that breaks rotational invariance has been also considered where gluon vortices are induced together with the spontaneous generation of a magnetic field Ferrer-Vortices . This, second scenario is motivated by the solution to the so-called ”zero-mode problem” that takes place for a charged spin-one field in the presence of a magnetic field larger than the field square Meissner mass Yang-Mills -String Theory .
@@
In the present work, we will show that in the presence of a magnetic field the unstable neutral 2SC phase will be stabilized by restructuring its ground state by the formation of a gluon vortex condensate as follows from the well-known solution of the ”zero-mode problem,” which has been already found in different contexts as for Yang-Mills fields in Yang-Mills , for the $W^{\pm}_{\mu}$ bosons of the electroweak theory in W-EW , for the charged gluons of the MCFL phase of CS in MCFL-Vortices and even for higher-spin fields in the context of string theory in String Theory . In the magnetized neutral 2SC system the gluon vortex solution will be even reinforced for certain density regions in the presence of a magnetic field as we will show in this work.
@@
Something else to notice is that an important characteristic of spin-zero CS phases is the lack of Meissner effect in contrast to the conventional superconductivity. Although the original electromagnetic $U(1)_{em}$ symmetry is broken by the formation of quark Cooper pairs, a residual $\widetilde{U}(1)$ symmetry still remains. The massless gauge field associated with this symmetry in the 2SC phase is given by the linear combination of the conventional photon field $A_{\mu}$ and the 8th-gluon field $G^{8}_{\mu}$ Rotated Fields as $\widetilde{A}_{\mu}=\cos\theta A_{\mu}-\sin\theta G^{8}_{\mu}$. The field $\widetilde{A}_{\mu}$ plays the role of a long-range in-medium electromagnetic field, which is also called the rotated electromagnetic field, while the orthogonal combination $\widetilde{G}_{\mu}^{8}=\sin\theta A_{\mu}+\cos\theta G^{8}_{\mu}$ is massive. The mixing angle depends on the coupling constants through the relation $\tan\theta=e/\sqrt{3}g$, with $g$ and $e$ the strong and electromagnetic coupling constants, respectively. In this paper we will investigate the effect of the rotated magnetic field associated with $\widetilde{A}_{\mu}$ on neutral 2SC matter and especially on the Meissner masses of gluons that acquire rotated electric charges, $G^{(4,5,6,7)}_{\mu}$, and which are the ones associated with the tachyonic modes in this medium. The mechanism that serves to remove in this context the chromomagnetic instability is due to the formation of gluonic vortices together with the generation of a magnetic field. Thus, we will also show how much the applied magnetic field can be boosted by the generation of the gluon vortices that can be induced at the moderately high baryon density expected to be reached in NS. Finally, based on those results, one of the main outcomes of this paper will be to present a natural mechanism to explain the possible origin of the strong magnetic fields exhibited by magnetars.
@@
The rest of the paper is organized as follows: In Section 2, we review the NJL formulation of the 2SC neutral phase of CS in the presence of a uniform magnetic field. The results for the gap, $\Delta$, and electric chemical potential, $\mu_{e}$, found in this section will be used then in the subsequent sections. In Section 3, the Meissner mass of the rotationally charged gluons is calculated for different values of the diquark coupling constant in the range of baryonic densities that are considered to be of interest for NS. It will be shown that for coupling constants not extremely high, the Meissner mass exhibits a tachyonic behavior at moderately high magnetic fields in the whole range of densities under consideration. This is what has been called the chromomagnetic instability phenomenon. It will be discussed how this instability is removed by redefining the system ground state through the formation of gluon vortices. A peculiar effect generated by these vortices is the amplification of the magnitude of the applied magnetic field. By numerical calculations we will see that small fields up to $10^{15}$ G will be boosted in 2 orders of magnitude. This effect will be used then in Section 4 as a mechanism to produce the high strength that the magnetar magnetic fields can reach in its interior. In Section 5 we present the concluding remarks.
@@ II Cooper pair condensate at ${\bm{B}}\neq 0$ in the neutral medium
We consider in this paper a two-flavor NJL model of massless quarks in the presence of an external constant and uniform rotated magnetic field (from now on we are going to ignore the term ”rotated” but it will be indicated by a wavy-hat notation on top of the magnitudes). The Lagrangian reads
${\cal L}=\bar{\psi}\left[i\gamma^{\mu}(\partial_{\mu}+i\tilde{e}\tilde{Q}\tilde{A}_{\mu})+\mu\gamma^{0}\right]\psi+G_{S}\left[(\bar{\psi}\psi)^{2}+(\bar{\psi}i\gamma^{5}{\tau^{3}}\psi)^{2}\right]+G_{D}\left(\bar{\psi}i\gamma^{5}T^{2}\tau^{2}\psi^{C}\right)\left(\bar{\psi}^{C}i\gamma^{5}T^{2}\tau^{2}\psi\right),$ (1)
where $\psi^{T}=(u,d)$ is the quark doublet and $\psi^{C}=C\bar{\psi}^{T}$ is its charge conjugate with $C=i\gamma^{2}\gamma^{0}$ the charge conjugation operator, and $G_{S}$ and $G_{D}$ the coupling constants of the scalar and diquark channels of the four-fermion interaction theory respectively. Here, $T^{2}$ is the second Gell-Mann matrix in color space and $\tau^{2,3}$ are Pauli matrices in flavor space. $\tilde{Q}$ is the rotated charge matrix defined as
@@
in units of $\tilde{e}=\frac{\sqrt{3}ge}{\sqrt{3g^{2}+e^{2}}}>0$. Here $Q_{f}={\rm diag}(2/3,-1/3)$ is the original electromagnetic charge matrix for quarks and $T_{c}^{8}={\rm diag}(1/\sqrt{3},1/\sqrt{3},-2/\sqrt{3})$ is the 8th Gell-Mann matrix. To be explicit, the $\tilde{Q}$ charges for different quarks are given in Table I and for different gluons in Table II.
@@
In this convention, the only diquark pairs are $u_{r}d_{g}$, and $u_{g}d_{r}$. These pairs are neutral with respect to $\tilde{Q}$. This is the reason why there is no Meissner effect for $\tilde{B}$ due to the absence of a charged ground state.
@@
The quark chemical potentials related with electrical and color neutralities, as well as $\beta$ equilibrium can be expressed by the following matrix in color-flavor space:
@@
The notations $\bar{\mu}=\mu-\frac{1}{6}\mu_{e}+\frac{1}{3}\mu_{8}$, $\delta\mu_{e}=\frac{1}{2}\mu_{e}$ and $\delta\mu_{8}=\mu_{8}$ have been used. Notice that with the definitions (5)-(8) the weak beta-equilibrium condition $\mu_{u}+\mu_{e}-\mu_{\nu_{e}}=\mu_{d}$ is satisfied for each quark color. The chemical potentials $\mu_{e}$ and $\mu_{8}$ are dynamical parameters that have to be determined from the electrical and color neutrality equations. The color chemical potential $\mu_{8}$ is usually very small compared with the other chemical potentials $\mu_{e}$ and $\mu$ mu-8 . We, therefore, will assume $\mu_{8}=0$ in the following calculations since as has been found by other authors it will not significantly change the results.
@@
which is obtained from Eq. (1) by assuming $\Delta=2G_{D}\langle{\bar{\psi}}^{C}i\gamma^{5}T^{2}\tau^{2}\psi\rangle$ and making the mean field approximation. Note that we are not considering the chiral condensation. From previous results Jaikumar-2 , we see that for the density region $\mu>320$ MeV the chiral condensate already vanished. Thus, in our calculations we will consider only the region with $\mu>320$ MeV, which on the other hand is the region where $\Delta\neq 0$ Jaikumar-2 .
@@
Figure 1: The solutions of the gap $\Delta$ and mismatch $\delta\mu_{e}$ versus baryonic chemical potential for different diquark couplings at a magnetic-field value of $10^{16}$.
Figure 2: The solutions of the gap $\Delta$ and mismatch $\delta\mu_{e}$ versus baryonic chemical potential for different diquark couplings at a magnetic-field value of $10^{18}$ G.
Thus, we obtain the thermodynamical potential
@@
with the contribution in the parentheses coming from the electrons, which must be introduced to maintain the system neutrality and $\beta$ equilibrium.
Carrying out the sum in Matsubara frequencies, we finally have
@@
In Figs. 1 and 2, we present for two different values of magnetic fields the solutions of the minimum equations (36) in the range of baryonic chemical potentials from 320 to 500 MeV, which are expected to correspond to the typical density region of compact stars. Comparing Figs. 1 and 2 we notice that the magnetic field does not significantly affect the behavior of $\Delta$ and $\delta\mu_{e}$ in the parameter range under consideration. As expected, the coupling strength $\eta$ affects significantly the value of the gap $\Delta$ producing a significant increase. The electric chemical potential, on the other hand, also increases with $\eta$. This is indicating that as more quarks having a positive conventional electric charge condense in Cooper pairs, the system needs a larger number of electrons to keep its electric neutrality.
Figure 3: The ratio of the squared Meissner mass of charged gluons to the magnetic mass for different diquark couplings at a magnetic-field value of $10^{16}$ G.
Figure 4: The ratio of the squared Meissner mass of charged gluons to the magnetic mass for different diquark couplings at a magnetic-field value of $10^{18}$ G.
@@
The gluon screening properties in the neutral 2SC phase can be studied by calculating the gluon polarization tensors considered in detail in Ref. Huang . In principle, for sufficiently high magnetic fields the polarization tensor of charged particles can change their Lorentz tensorial structure showing a space anisotropy determined by the field direction Pi_Anisotropy . Nevertheless, for astrophysical applications to compact objects since $eB\ll\mu^{2}$, the anisotropy in the Lorentz structure of the polarization operator becomes a second-order effect that can be neglected, and, hence we can keep the polarization operator structure characteristic of a dense medium at zero field reducing the effect of the magnetic field to its dependence on $\Delta$ and $\mu_{e}$.
In Huang it was found that while the Meissner masses of the three gluons that correspond to the unbroken $SU(2)_{c}$ subgroup are vanishing as expected, the charged gluons that correspond to the broken generators of the $SU(3)_{c}$ group could be imaginary in some parameter regions. The explicit Meissner masses for the charged and the rotated $8$th gluon fields are given, respectively, by